OpenStudy (anonymous):

help i did this but it saids my answer is incorrect any help thanks: ((a4)/(2a2-4a4))*((8a3+1)/(2a+4a2)) Final result : -a • (4a2 - 2a + 1) ——————————————————— 4 • (2a2 - 1)

OpenStudy (anonymous):

is that a^4

OpenStudy (anonymous):

Is this your original form? \[ \frac{a^4 \left(8 a^3+1\right)}{\left(4 a^2+2 a\right) \left(2 a^2-4 a^4\right)} \]

OpenStudy (anonymous):

just use this formula\[a ^{3}+b ^{3}=(a+b)(a ^{2}+b ^{2}-ab)\]

OpenStudy (anonymous):

May be they expect you to put your answer like what I found \[ \frac{a^4 \left(8 a^3+1\right)}{\left(4 a^2+2 a\right) \left(2 a^2-4 a^4\right)}=\frac{a \left(4 a^2-2 a+1\right)}{4-8 a^2} \]

OpenStudy (anonymous):

Your answer is correct. It is the same as mine.